# SliM-LLM: Salience-Driven Mixed-Precision Quantization for Large Language Models

### v1

@@ F.2 Distribution of salience, activation and weight magnitude
In line with previous findings [32, 44], activations demonstrate particularly marked outlier phenomena on anomalous tokens and channels, with extremes differing by more than two orders of magnitude. Notably, distinct anomalous channels are present in the MHA’s Query, Key, and Value layers, where outliers vary significantly across different tokens. This pattern is consistent in the FFB layers. We observe that disparities in weight magnitudes are less pronounced than those in activation, thus exerting a reduced impact on outlier channels. Moreover, weights exhibit structured distributions along rows or columns [12, 20], affecting the overall distribution of salience from a row-wise perspective (Fig. 9). However, the most prominent salience is predominantly driven by activation across channels (column-wise).
F.3 Hessian Diagonal Clustering
Sec. 3.2.1 demonstrates that outlier tokens in input activations result in significant values at the corresponding positions along the diagonal of the weight Hessian matrix. Additionally, due to the token sink phenomenon [45, 32], areas around significantly activated key tokens exhibit increased salience, creating clusters of salient regions along the Hessian matrix diagonal. To further elucidate this phenomenon, Fig. 10 shows the values along the diagonal of the Hessian matrix for selected weights in the $2^{nd}$ and $10^{th}$ layers of the LLaMA-7B model. Within this diagonal, certain positions display pronounced values (indicated in red), whereas others are relatively moderate. In the attention aggregation layer of the $10^{th}$ layer, the token sink phenomenon results in a pronounced convergence of significant values along the Hessian matrix diagonal, with deep red areas indicating regional clustering. These findings reinforce the influence of input activations on the diagonal of the Hessian matrix, subsequently leading to a clustering phenomenon in the salience distribution of weights across channels.
@@ F.3 Hessian Diagonal Clustering
Appendix G More Comparisons
In this section, we provide supplementary experiments for SliM-LLM. Tab. 6 displays the comparative results of SliM-LLM and SliM-LLM* with other methods on the OPT series models. Tab. 7 shows the performance of SliM-LLM when quantizing the LLaMA family models on the C4 dataset, while Tab. 8 also compares the results of SliM-LLM* on the C4 dataset.
@@ Appendix G More Comparisons
Appendix H Real Dialog Examples

### v2

@@ G.2 Distribution of salience, activation and weight magnitude
In line with previous findings (Nrusimha et al., 2024; Xiao et al., 2023a), activations demonstrate particularly marked outlier phenomena on anomalous tokens and channels, with extremes differing by more than two orders of magnitude. Notably, distinct anomalous channels are present in the MHA’s Query, Key, and Value layers, where outliers vary significantly across different tokens. This pattern is consistent in the FFB layers. We observe that disparities in weight magnitudes are less pronounced than those in activation, thus exerting a reduced impact on outlier channels. Moreover, weights distribute structurally along rows or columns (Dettmers et al., 2023; Huang et al., 2024a), affecting the overall distribution of salience from a row-wise perspective (Fig. 9). However, the most prominent salience is predominantly driven by activation across channels (column-wise).
G.3 Hessian Diagonal Clustering
Sec. 3.2.1 demonstrates that outlier tokens in input activations result in significant values at the corresponding positions along the diagonal of the weight Hessian matrix. Additionally, due to the token sink phenomenon (Xiao et al., 2023b; Nrusimha et al., 2024), areas around significantly activated key tokens exhibit increased salience, creating clusters of salient regions along the Hessian matrix diagonal. To further elucidate this phenomenon, Fig. 10 shows the values along the diagonal of the Hessian matrix for selected weights in the $2^{nd}$ and $10^{th}$ layers of the LLaMA-7B model. Within this diagonal, certain positions display pronounced values (indicated in red), whereas others are relatively moderate. In the attention aggregation layer of the $10^{th}$ layer, the token sink phenomenon results in a pronounced convergence of significant values along the Hessian matrix diagonal, with deep red areas indicating regional clustering. These findings reinforce the influence of input activations on the diagonal of the Hessian matrix, subsequently leading to a clustering phenomenon in the salience distribution of weights across channels.
@@ G.3 Hessian Diagonal Clustering
Appendix H More Comparisons
In this section, we provide supplementary experiments for SliM-LLM. Tab. 8 displays the comparative results of SliM-LLM and SliM-LLM+ with other methods on the OPT series models. Tab. 9 shows the performance of SliM-LLM when quantizing the LLaMA family models on the C4 dataset, while Tab. 10 also compares the results of SliM-LLM+ on the C4 dataset. In Tab. 11, we compared the quantization results of GPTQ, AWQ, and SliM-LLM at 2-bit on the Gemma2 and Mixtral models, demonstrating the greater stability of SliM-LLM across a wider range of model structures. Additionally, in Tab. 12, we supplemented the 4-bit results of different quantization methods in the LLaMA series models, showing that SliM-LLM and SliM-LLM+ exhibit the smallest quantization errors at practical 4-bit levels. To provide a comprehensive evaluation across a broader set of benchmarks, we further compared the quantization results on MMLU and MathQA in Tab. 13.
Appendix I Real Dialog Examples
Appendix J Model Experiments on Efficiency
Tab. 14 presents further efficiency comparisons between AutoGPTQ and SliM-LLM. The results shows that even under the larger LLMs with 70B parameters, SliM-LLM can also show the better accuracy with comparable inference efficiency.
@@ Appendix I Real Dialog Examples
Table 11: PPL Comparison on Gemma2 and Mixtral.
Model/Evaluation Method PPL (wikitext2)
Gemma2-9B GPTQ 2-bit 186.77
AWQ 2-bit 217.83
SliM-LLM 2bit 26.30
Mixtral 8x7B GPTQ 2-bit 16.38
AWQ 2-bit 3.2e5
SliM-LLM 2bit 7.44
Table 12: The PPL results of our proposed method and other methods under 4bit quantization.
Method LLaMA-7B LLaMA-13B LLaMA2-7B LLaMA2-13B LLaMA3-8B
FP16 5.68 5.09 5.47 4.88 5.75
AWQ 5.81 5.30 5.62 4.97 6.63
GPTQ 5.85 5.20 5.61 4.98 6.50
SliM-LLM 5.83 5.16 5.59 4.95 6.42
Omniquant 5.77 - 5.58 - -
SliM-LLM+ 5.75 - 5.57 - -
Appendix K Quantization Performance on Vision Language Models
To further showcase the potential application capabilities of SliM-LLM, in Tab. 15, we deploy SliM-LLM on LLaVA-Next-8B (Liu et al., 2024) evaluated on 4 benchmarks. The results show that GPTQ, AWQ and SliM-LLM show comparable performance under the 3-bit context. However, when the bit-width is setting as 2, GPTQ and AWQ failed to generate the reasonable answer in each benchmark and get "N" results. SliM-LLM successfully generate the reasonable output and the accuracy is closed to 3-bit model, which presents the superior usability of SliM-LLM to wider environments.
Table 13: The results(%) on MMLU and MathQA for multiple quantized LLaMA models.
Model Method Humanities Social Sciences STEM Other MMLU MathQA
LLaMA-7B GPTQ 2-bit 24.87 21.84 21.79 24.01 23.32 21.11
AWQ 2-bit 24.21 21.71 21.25 23.98 22.95 22.21
SliM-LLM 2bit 24.94 23.60 23.40 25.50 25.10 23.74
LLaMA-13B GPTQ 2-bit 24.23 23.20 22.99 24.78 23.85 21.68
AWQ 2-bit 24.17 31.07 28.61 25.14 26.89 21.98
SliM-LLM 2bit 25.12 31.74 29.19 26.17 27.05 23.17
LLaMA2-7B GPTQ 2-bit 25.02 22.13 22.61 23.17 23.44 21.07
AWQ 2-bit 25.12 22.79 24.26 24.01 24.51 19.06
SliM-LLM 2bit 26.60 23.23 25.70 25.70 25.81 22.55
LLaMA2-13B GPTQ 2-bit 23.91 27.17 26.10 25.78 25.53 20.87
AWQ 2-bit 24.17 31.07 28.61 25.14 26.89 19.53
SliM-LLM 2bit 26.27 32.20 29.98 26.46 27.34 23.48
Table 14: LLaMA-2-70B results of GPTQ and Slim-LLM on GPU. Group size is set to 128.
[…]
3-bit GPTQ 28.0G 34.9G 3.85 6.5
SliM-LLM 28.0G 35.2G 3.67 6.2
2-bit GPTQ 16.4G 23.3G 8.78 9.7
SliM-LLM 23.5G 4.4G 6.28 8.4
Table 15: Quantization results on Vision Language Models (VLMs).
#W #G AI2D ChartQA DocVQA MMBench
GPTQ 3 128 66.2 65.1 75.6 67.4
2 128 N N N N
AWQ 3 128 67.7 65.4 74.4 68.0
2 128 N N N N
SliM-LLM 3 128 68.2 67.5 74.8 68.9
2 128 57.2 49.3 60.6 60.9